OpenStudy (summersnow8):

How many Newtons of force are required to stop a truck 2250 kg truck going 32.0 km/h in 4.00s? @rajat97

OpenStudy (shamim):

I m not solving ur problem. Because u did not give me medal in ur previous problem

OpenStudy (rajat97):

here, we just need to find the acceleration of the truck(i mean deceleration) and then multiply it with the mass of the truck first we convert everything into SI units the initial speed=u=32km/hr=8.89 m/s the final speed=v=0m/s time=t=4s we can use the kinematical equation v=u+at where a is the acceleration and all the other variables are defined above so from here, you get the acceleration (or the deceleration as we want to stop the truck) of the truck and then we just need to multiply the value of a i.e. acceleration with the mass of the truck i.e. 2250Kg

OpenStudy (rajat97):

i may not be very clear so just feel free to ask anything

OpenStudy (summersnow8):

how do you find Newtons from all this information?

OpenStudy (rajat97):

by Newtons in the question, they mean 'Force' and we know, force is F=mass x ?

OpenStudy (summersnow8):

\[a = \frac{ V-u }{ t } = \frac{ 0-8,888 }{ 4.00 } = -2.22\] \[F = m * a = (2250 * -2.222) = -5000\]

OpenStudy (summersnow8):

*8.888

OpenStudy (rajat97):

Awesome! you are absolutely right!

OpenStudy (summersnow8):

do i keep the - sign?

OpenStudy (rajat97):

the -ve sign just means that the force is applied in opposite direction to that of the motion of the object which is the truck here

OpenStudy (summersnow8):

so if i submit an answer, should I submit 5000 or -50000

OpenStudy (rajat97):

you should submit just 5000 as they have asked 'how many newtons' and not the direction of the force i.e. only magnitude

OpenStudy (summersnow8):

okay thanks for your help

OpenStudy (rajat97):

any time girl!